OpenStudy (anonymous):

Find the vertex of the function y = -x^2 + x Question 8 options: 1) (1/2, 1/4) 2) (1/4, 1/2) 3) (1/2, 1/2) 4) (1/4, 1/4)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1

OpenStudy (triciaal):

for an equation in general form ax + by + c = 0 the x coordinate of the vertex is -b/2a and the y value will be the result when that x value is plugged into the function

OpenStudy (anonymous):

x=-b/(2a)=-1(2*-1)=1/2 so plug into the original equation and get y=1/4

OpenStudy (anonymous):

thank you!

OpenStudy (triciaal):

you are welcome!